The volume of a gas is 400.0 mL when the pressure is 1.00 atm. At the same temperature, what is the pressure at which the volume of the gas is 2.0L?
1
Expert's answer
2013-03-11T10:08:56-0400
P1V1/T1 = P2V2/T2
If one of these values is not given then it is considered a constant and you just leave that variable out of the equation.
PV=nRT is used when you are working the # of moles or grams. n is the # of moles so if given grams you must convert to moles before you begin.
STP (standard temperature and pressure) P= 1atm ,T= 273K, R= 0.08206 L-atm / mol-K
P1V1= P2V2

P2 =P1V1/V2 =1 *0.4 / 2 = 0.2 atm